OpenStudy (henrietepurina):

Which numbers should be multiplied to obtain 165^2 - 123^2?

OpenStudy (henrietepurina):

@paki @TheSmartOne

jagr2713 (jagr2713):

is there an answer choice

jagr2713 (jagr2713):

@HenrietePurina

OpenStudy (henrietepurina):

yes there are @jagr2713

OpenStudy (henrietepurina):

42 and 42 42 and 288 1,764 and 1,764 4,202 and 4,248

OpenStudy (henrietepurina):

sorry for the wait, I went to get some water :P

OpenStudy (henrietepurina):

are you understanding this question @jagr2713 ?

jagr2713 (jagr2713):

yes give me a second

OpenStudy (henrietepurina):

kk :D

jagr2713 (jagr2713):

a2−b2=(a−b)(a+b)

OpenStudy (henrietepurina):

ok.... lets try putting the variables in :D 1652−1232=(165−123)(165+123)

jagr2713 (jagr2713):

yep so whats the answer?

OpenStudy (henrietepurina):

oh gosh sorry 165^2−123^2=(165−123)(165+123)

jagr2713 (jagr2713):

what do you get when you do 165+123?

OpenStudy (henrietepurina):

12096 that is for both

OpenStudy (henrietepurina):

sorry, didn't see your question :D

jagr2713 (jagr2713):

what do you get?

OpenStudy (henrietepurina):

12096

OpenStudy (henrietepurina):

or true, that yes they are equal to each other

jagr2713 (jagr2713):

its not 12096 what is 165 plus 123

OpenStudy (henrietepurina):

288

jagr2713 (jagr2713):

yep and what is 165-123?

OpenStudy (henrietepurina):

42

jagr2713 (jagr2713):

so whats the answer?

OpenStudy (henrietepurina):

288*42=12096

jagr2713 (jagr2713):

it is not an answer choice D;

jagr2713 (jagr2713):

it is B :D

OpenStudy (henrietepurina):

yeah, I though about that by thinking what numbers multiply to make 12096 :P

jagr2713 (jagr2713):

ok but the answer is B:D

OpenStudy (henrietepurina):

yeah... thanks btw :D

jagr2713 (jagr2713):

yw :D